# Exponential Sums with Sparse Polynomials and Distribution of the Power Generator

### v1

@@ Abstract.
We obtain new bounds on complete rational exponential sums with sparse polynomials modulo a prime, under some mild condition on the degrees of the monomials of such polynomials. These bounds, when they apply, give explicit versions of a result of J. Bourgain (2005). We then use one of these bounds to study the multidimensional distribution of the classical power generator of pseudorandom numbers, which has not been possible within previously known results.
@@ 1.1. Set-up and motivation
Let $p$ be a prime. Given a polynomial $f\in{\mathbb{F}}_{p}[X]$ over a finite field ${\mathbb{F}}_{p}$ of $p$ elements, where $p$ is prime, we define the Weil sum as
(1.1) $S(f)=\sum_{x\in{\mathbb{F}}_{p}}{\mathbf{\,e}}_{p}(f(x)),$
where
${\mathbf{\,e}}_{p}(z)=\exp(2\pi iz/p),$
and we always assume that the elements of ${\mathbb{F}}_{p}$ are represented by the set $\{0,\ldots,p-1\}$.
The celebrated result of Weil [38] asserts that for any non-constant polynomial $f\in{\mathbb{F}}_{p}[X]$, we have
(1.2) $\left|S(f)\right|\leq(n-1)p^{1/2},$
where $n=\deg f$, see also, for example, [19, Chapter 11] and [23, Chapter 6].
Clearly the bound (1.2) becomes trivial for $n\geq p^{1/2}+1$, which is the limit of our present knowledge. However, in the case of so-called sparse polynomials, that is, polynomials of the form
(1.3) $f(x)=\sum_{i=1}^{r}a_{i}x^{e_{i}}\in{\mathbb{F}}_{p}[X]$
with a small (compared to the degree) number $r$ of distinct exponents $p-1\geq e_{r}>\ldots>e_{1}\geq 1$, and $a_{1},\ldots,a_{r}\in{\mathbb{F}}_{p}^{*}$, better bounds are know.
Most of attention was dedicated to the monomial case, that is, to $r=1$. Since the first improvement of the general bound (1.2) given in [32], several stronger bounds have been obtained, see, for example [18, 21, 30] and references therein. We observe, that if $r=1$, then without loss of generality, we can assume that $d=e_{1}\mid p-1$, which reduces the corresponding sums to estimating exponential sums of elements of subgroups of ${\mathbb{F}}_{p}^{*}$ of order $(p-1)/e$. In particular, by a result of Shkredov [30, Theorem 1], for $f(X)=aX^{d}$ with $a\in{\mathbb{F}}_{p}^{*}$, we have
(1.4) $S(f)\ll d^{1/2}p^{2/3}(\log p)^{1/6},$
@@
For $r=2$ a large variety of bounds is available [17, 26, 27, 34, 35], which require rather strong restrictions on some greatest common divisors associated with the powers $d$ and $e$ of monomials in $f(X)=aX^{d}+bX^{e}\in{\mathbb{F}}_{p}[X]$. For our purpose we use a result of Cochrane and Pinner [8, Theorem 1.3]:
(1.5) $|S(f)|\leq\gcd(d-e,p-1)+2.292g^{13/46}p^{89/92}$
@@
(1.6) $\left|S(f)\right|\leq p^{1-\delta},$
with some $\delta>0$ depending only on $r$ and $\varepsilon>0$. Unfortunately, Bourgain [3] has not provided any explicit formulas for $\delta$ as a function of $r$ and $\varepsilon>0$, while an attempt to do so in [28, Theorems 4 and 5] seems to be faulty. In particular, it appears that the argument in [28] misquotes the result of [31, Corollary 16], which, after correcting, leads to exponentially smaller saving (and yet the rest of [28] is still to be verified).
We also recall than the bound (1.5) has been used in [25] to study the uniformity of distribution properties of the power generator of pseudorandom numbers, More precisely, given an integer $e\geq 2$ and a prime $p$, this generator outputs the sequence $\{u_{n}\}$, defined by
@@
1.2. Exponential sums with sparse polynomials
Here we use the bounds (1.4) and (1.5) as the basis for the inductive argument of Bourgain [3], in the form outlined by Garaev [16, Section 4.4] to derive a bound on exponential sums $S(f)$ with sparse polynomials $f$ as in (1.3).
@@ 1.2. Exponential sums with sparse polynomials
Using Theorems 1.1 and 1.3 combined with the Koksma–Szüsz inequality, see Lemma 3.1, one can get explicit versions of a result of Bombieri, Bourgain and Konyagin [2, Theorem 17], see also [5, Corollary 9], counting the number of solutions to the system of congruences
@@
There are also bounds on exponential sums with sparse polynomials modulo composite numbers [4, 33], which are stronger than a generic bound for dense polynomials. We now recall that the approach of Bourgain [4] makes use of the bound (1.6). Hence, it is natural to ask whether Theorems 1.1 and 1.3 can lead to appropriate explicit versions of  [4, Theorem 1].
@@ 1.3. Multidimensional distribution of the power generator
For $s=1$, a bound on $D_{e,\vartheta,1}(p,N)$ is given by [25, Corollary 3]. However, the approach of [25] does not extend to $s\geq 2$. Here we use Theorem 1.3 to establish such a result.
Clearly, if $\gcd(e,p-1)=1$ and $\gcd(\vartheta,p)=1$, then the sequence (1.7) is purely periodic.
@@ 2.1. Preliminaries
We use the inductive idea of Bourgain [3], however we apply it in the form outlined by Garaev [16, Section 4.4], also applied in [1, 27]. It is important to observe that the basis of induction is formed by the cases $r=1$ and $r=2$ (just $r=1$ is not enough).
We prove both Theorems 1.1 and 1.3 simultaneously, giving more details for the proof of Theorem 1.1, which is more technically involved.
Let us consider the exponential sum $S(f)$ as in (1.1) with polynomials $f$ as in (1.3). We also use $I_{r,t}$ to denote the number of solutions to the system of equations
@@
$S(f)^{t}$ $=\sum_{x_{1},\ldots,x_{t}\in{\mathbb{F}}_{p}}{\mathbf{\,e}}_{p}\left(f\left(x_{1}\right)+\ldots+f\left(x_{t}\right)\right)$
@@
$|S(f)|^{t}$ $\ll p^{-1}\sum_{\lambda_{1},\ldots,\lambda_{r}\in{\mathbb{F}}_{p}}J_{r,t}(\lambda_{1},\cdots,\lambda_{r})\left|\sum_{z\in{\mathbb{F}}_{p}^{*}}\mathbf{e}_{p}\left(\lambda_{1}z^{e_{1}}+\cdots+\lambda_{r}z^{e_{r}}\right)\right|.$
@@
$\left|S(f)\right|^{2st}$ $\ll p^{-2s}\left(\sum_{\lambda_{1},\ldots,\lambda_{r}\in{\mathbb{F}}_{p}}J_{r,t}(\lambda_{1},\cdots,\lambda_{r})\right)^{2s-2}$
@@
where $\mathbf{0}$ is a zero vector in ${\mathbb{F}}_{p}^{r-1}$. Thus we estimate $I_{r,t}$ via exponential sums with $(r-1)$-sparse polynomials, which enable inductive approach. Furthermore, for $t=2$ we use the following bound which is a weakened form or a result of Cochrane and Pinner [8, Theorem 7.1].
@@
Let $d=\gcd(e_{1},e_{2},p-1)$. Assume that $p$ is sufficiently large. By [8, Theorem 7.1] we have the bound
@@
Using the same argument as in the proof of Lemma 2.2, while taking $d=1$, we also see that [8, Theorem 7.1] leads to the following simpler version of Lemma 2.2.
@@ 2.2. Induction basis
First we observe that the bounds (1.4) and (1.5) ensure that under the condition of Theorems 1.1 and 1.3 the desired results hold for $r=1$ and $r=2$.
@@
$S(f)\ll d^{1/2}p^{2/3}(\log p)^{1/6}\ll\left(p^{7/26+14\varepsilon/13}\right)^{1/2}p^{2/3}(\log p)^{1/6}\ll p^{1-\varepsilon},$
@@
Furthermore, under the assumption of $e_{1}$ in Theorem 1.3 we obviously have $S(f)=0$. Hence, we can take $\kappa_{1}(\varepsilon)=\varepsilon$.
@@
$S(f)\ll p^{8/13+32\varepsilon/13}+\left(p^{3/26-46\varepsilon/13}\right)^{13/46}p^{89/92}\ll p^{1-\varepsilon}$
@@
$S(f)\ll p^{16/23}+p^{89/92}\leq p^{89/92}=p^{1-3/92}.$
Hence, recalling the definition (1.8), we see that this establishes the basis of induction for Theorems 1.1 and 1.3.
@@ 2.3. Induction step
where $t_{r}(\varepsilon)$ is given by (1.10), to ensure that (2.5) holds, we derive $S(f)\ll p^{1-\varepsilon/t_{r}(\varepsilon)}=p^{1-\kappa_{r}(\varepsilon)}$, where $\kappa_{r}(\varepsilon)$ is as in (1.9).

### v2

@@ Abstract.
We obtain new bounds on complete rational exponential sums with sparse polynomials modulo a prime, under some mild conditions on the degrees of the monomials of such polynomials. These bounds, when they apply, give explicit versions of a result of J. Bourgain (2005). In turn, as an application, we also obtain an explicit version of a result of J. Bourgain (2010) on rational exponential sums with sparse polynomials modulo an arbitrary composite number. We then use one of these bounds to study the multidimensional distribution of the classical power generator of pseudorandom numbers, which has not been possible within previously known results.
@@ 1.1. Set-up and motivation
Let $q$ be a positive integer. Given a polynomial $f\in{\mathbb{Z}}_{q}[X]$ over the residue ring we define the following exponential sums
(1.1) $S_{q}(f)=\sum_{x\in{\mathbb{Z}}_{q}^{*}}{\mathbf{\,e}}_{q}(f(x)),$
over the group of units ${\mathbb{Z}}_{q}^{*}$, where
${\mathbf{\,e}}_{q}(z)=\exp(2\pi iz/q),$
and we always assume that the elements of ${\mathbb{Z}}_{q}$ are represented by the set $\{0,\ldots,q-1\}$.
Here we are mostly interested in sparse polynomials, that is, polynomials of the form
(1.2) $f(X)=\sum_{i=1}^{r}a_{i}X^{e_{i}}\in{\mathbb{Z}}_{q}[X],$
with a small (compared to the degree) number $r$ of distinct exponents $e_{r}>\ldots>e_{1}\geq 1$, for which we always assume
(1.3) $\gcd(a_{1},\ldots,a_{r},q)=1$
We first discuss the case of prime $q=p$, where we use ${\mathbb{F}}_{p}={\mathbb{Z}}_{p}$ to denote the finite field of $p$ elements.
The celebrated result of Weil [43] asserts that for any non-constant polynomial $f\in{\mathbb{F}}_{p}[X]$, over a finite field ${\mathbb{F}}_{p}={\mathbb{Z}}_{p}$ of $p$ elements, where $p$ is prime, we have
(1.4) $\left|S_{p}(f)\right|\leq(d-1)p^{1/2},$
where $d=\deg f$, see also, for example, [21, Chapter 11] and [25, Chapter 6].
Clearly the bound (1.4) becomes trivial for $d\geq p^{1/2}+1$, which is the limit of our present knowledge. However, in the case of sparse polynomials (1.2) where we can also assume $p-1\geq e_{r}>\ldots>e_{1}\geq 1$ and $a_{1},\ldots,a_{r}\in{\mathbb{F}}_{p}^{*}$, better bounds are known.
Most of attention was dedicated to the monomial case, that is, to $r=1$. Since the first improvement of the general bound (1.4) given in [35], several stronger bounds have been obtained, see, for example [20, 23, 33] and references therein. We observe, that if $r=1$, then without loss of generality, we can assume that $d=e_{1}\mid p-1$, which reduces the corresponding sums to estimating exponential sums of elements of subgroups of ${\mathbb{F}}_{p}^{*}$ of order $(p-1)/e$. In particular, by a result of Shkredov [33, Theorem 1], for $f(X)=aX^{d}$ with $a\in{\mathbb{F}}_{p}^{*}$, we have
(1.5) $S_{p}(f)\ll d^{1/2}p^{2/3}(\log p)^{1/6},$
@@
For $r=2$ a large variety of bounds is available [19, 28, 29, 37, 38], which require rather strong restrictions on some greatest common divisors associated with the powers $d$ and $e$ of monomials in $f(X)=aX^{d}+bX^{e}\in{\mathbb{F}}_{p}[X]$. For our purpose we use a result of Cochrane and Pinner [9, Theorem 1.3]:
(1.6) $|S_{p}(f)|\leq\gcd(d-e,p-1)+2.292g^{13/46}p^{89/92}$
@@
(1.7) $\left|S_{p}(f)\right|\leq p^{1-\delta},$
with some $\delta>0$ depending only on $r$ and $\varepsilon>0$. Unfortunately, Bourgain [3] has not provided any explicit formulas for $\delta$ as a function of $r$ and $\varepsilon>0$, while an attempt to do so in [30, Theorems 4 and 5] seems to be faulty. In particular, it appears that the argument in [30] misquotes the result of [34, Corollary 16], which, after correcting, leads to exponentially smaller saving (and yet the rest of [30] is still to be verified).
For composite $q$ and “dense” polynomials $f(X)=a_{d}X^{d}+\ldots+a_{1}X\in{\mathbb{Z}}_{q}[X]$ of degree $d$ with $\gcd(a_{1},\ldots,a_{r},q)=1$, we have
$S_{q}(f)\ll q^{1-1/d},$
see, for example, [10, 39]. Note that this is a slight improvement of the classical bound of Hua (containing an additional factor $q^{o(1)}$), see, for example, [42, Theorem 7.1]. For sparse polynomials as in (1.2) and (1.3), but of fixed degree $\deg f=e_{r}$ a better bound of the type
(1.8) $|S_{q}(f)|\leq q^{1-1/r+o(1)}$
is given in [36, Theorem 1].
Bourgain [4], using the bound (1.7), has given several estimates for the sums $S_{q}(f)$ with sparse polynomials as in (1.2) and (1.3), For example, by [4, Theorem 1], for any $\eta>0$ there exist $\delta>0$ depending only on $r$ and $\eta>0$, such that for
(1.9) $\left|S_{q}(f)\right|\leq q^{1-\delta},$
provided that
$q\geq\exp\left(d^{\eta}\right),$
where $d=\deg f=e_{r}$. In particular, the derivation of (1.9) in [4] is based on (1.7) from [4]. Here, despite that our explicit version of (1.7) is more restrictive in terms of the conditions on the exponents $e_{1},\ldots,e_{r}$, it is sufficient to get a version of (1.9) with an explicit value of $\delta$. A further improvement comes from the new way we treat exponential sums modulo $p^{m}$, when $m\geq 2$ is not too large, see Lemma 4.2 below.
1.2. Applications
We also recall that the bound (1.6) has been used in [27] to study the uniformity of distribution properties of the power generator of pseudorandom numbers, More precisely, given an integer $e\geq 2$ and a prime $p$, this generator outputs the sequence $\{u_{n}\}$, defined by
@@ 1.2. Applications
2. Main results
2.1. Exponential sums with sparse polynomials over ${\mathbb{F}}_{p}$
Here we use the bounds (1.5) and (1.6) as the basis for the inductive argument of Bourgain [3], in the form outlined by Garaev [18, Section 4.4] to derive a bound on exponential sums $S_{p}(f)$ with sparse polynomials $f$ as in (1.2).
@@ 2.1. Exponential sums with sparse polynomials over ${\mathbb{F}}_{p}$
Using Theorems 2.1 and 2.3 combined with the Koksma–Szüsz inequality, see Lemma 5.1, one can get explicit versions of a result of Bombieri, Bourgain and Konyagin [2, Theorem 17], see also [5, Corollary 9], counting the number of solutions to the system of congruences
@@
2.2. Exponential sums with sparse polynomials over ${\mathbb{Z}}_{q}$
We merge the bound of Theorem 2.3 with the argument Bourgain [4] to derive the following explicit version of (1.9) .
We now define
(2.5) $\sigma_{r}=\min\{\kappa_{r}(3/184),1/(50r^{3})\}.$
In fact $\sigma_{r}<\kappa_{r}(3/184)$ only for $r=2$, that is, alternatively we can define
$\sigma_{r}=\begin{cases}\kappa_{r}(3/184),&\text{for}\ r=1\leavevmode\nobreak\ \mathrm{and}\leavevmode\nobreak\ r\geq 3,\\ 1/400,&\text{for}\ r=2.\end{cases}$
Theorem 2.5.
Let $q\geq 2$ be an integer and let $f$ be as  (1.2) and (1.3) of degree $d$. Then uniformly over $d$, we have
$S_{q}(f)\ll\exp\left(d^{18}\right)q^{1-\sigma_{r}+o(1)},\qquad\text{as}\ q\to\infty,$
where $\sigma_{r}$ is given by (2.5).
@@ 2.2. Exponential sums with sparse polynomials over ${\mathbb{Z}}_{q}$
We emphasise that the main point of Theorem 2.5 is that the degree $d$ may grow with $q$, as the saving $\sigma_{r}$ and the implied constant depend only on the sparsity $r$ rather than on $d$, while the rate of decay of $o(1)$ in the exponent of $q$ depends only on $q$. Clearly, for a fixed $d$, the bound (1.8) of [36, Theorem 1] is much stronger.
Remark 2.7.
The power $18$ of $d$ in Theorem 2.5 is not optimised and can easily be reduced and the cost of redefining and reducing $\sigma_{r}$ as well. We have also ignored the factor $1/50$ at the front of $d^{18}$ which our argument gives, see the final estimate in the proof of Theorem 2.5.
Remark 2.8.
The limit of our technique with respect to the dependence on $d$ is $d^{26/3+\varepsilon}$, with any $\varepsilon>0$, instead of $d^{18}$ in Theorem 2.5 due to (4.3c). However, if we impose that the exponents $e_{i}$ are pairwise coprime, this issue disappears, leaving (4.3a) as the bottleneck. Thus, tightening (4.3a) and (4.3b) according to the conditions in Theorem 2.1, provides a much smaller exponent than $26/3$.
@@ 2.3. Multidimensional distribution of the power generator
For $s=1$, a bound on $D_{e,\vartheta,1}(p,N)$ is given by [27, Corollary 3]. However, the approach of [27] does not extend to $s\geq 2$. Here we use Theorem 2.3 to establish such a result.
Clearly, if $\gcd(e,p-1)=1$ and $\gcd(\vartheta,p)=1$, then it is easy to see that the sequence (1.10) is purely periodic.
@@ 3.1. Preliminaries
We use the inductive idea of Bourgain [3], however we apply it in the form outlined by Garaev [18, Section 4.4], also applied in [1, 29]. It is important to observe that the basis of induction is formed by the cases $r=1$ and $r=2$ (just $r=1$ is not enough).
We prove both Theorems 2.1 and 2.3 simultaneously, giving more details for the proof of Theorem 2.1, which is more technically involved.
Let us consider the exponential sum $S_{p}(f)$ as in (1.1) with polynomials $f$ as in (1.2). We also use $I_{r,t}$ to denote the number of solutions to the system of equations
@@
$S_{p}(f)^{t}$ $=\sum_{x_{1},\ldots,x_{t}\in{\mathbb{F}}_{p}}{\mathbf{\,e}}_{p}\left(f\left(x_{1}\right)+\ldots+f\left(x_{t}\right)\right)$
@@
$|S_{p}(f)|^{t}$ $\ll p^{-1}\sum_{\lambda_{1},\ldots,\lambda_{r}\in{\mathbb{F}}_{p}}J_{r,t}(\lambda_{1},\cdots,\lambda_{r})\left|\sum_{z\in{\mathbb{F}}_{p}^{*}}\mathbf{e}_{p}\left(\lambda_{1}z^{e_{1}}+\cdots+\lambda_{r}z^{e_{r}}\right)\right|.$
@@
$\left|S_{p}(f)\right|^{2st}$ $\ll p^{-2s}\left(\sum_{\lambda_{1},\ldots,\lambda_{r}\in{\mathbb{F}}_{p}}J_{r,t}(\lambda_{1},\cdots,\lambda_{r})\right)^{2s-2}$
@@
where $\mathbf{0}$ is a zero vector in ${\mathbb{F}}_{p}^{r-1}$. Thus we estimate $I_{r,t}$ via exponential sums with $(r-1)$-sparse polynomials, which enable inductive approach. Furthermore, for $t=2$ we use the following bound which is a weakened form or a result of Cochrane and Pinner [9, Theorem 7.1].
@@
Let $d=\gcd(e_{1},e_{2},p-1)$. Assume that $p$ is sufficiently large. By [9, Theorem 7.1] we have the bound
@@
Using the same argument as in the proof of Lemma 3.2, while taking $d=1$, we also see that [9, Theorem 7.1] leads to the following simpler version of Lemma 3.2.
@@ 3.2. Induction basis
First we observe that the bounds (1.5) and (1.6) ensure that under the condition of Theorems 2.1 and 2.3 the desired results hold for $r=1$ and $r=2$.
@@
$S_{p}(f)\ll d^{1/2}p^{2/3}(\log p)^{1/6}\ll\left(p^{7/26+14\varepsilon/13}\right)^{1/2}p^{2/3}(\log p)^{1/6}\ll p^{1-\varepsilon},$
@@
Furthermore, under the assumption of $e_{1}$ in Theorem 2.3 we obviously have $S_{p}(f)=0$. Hence, we can take $\kappa_{1}(\varepsilon)=\varepsilon$.
@@
$S_{p}(f)\ll p^{8/13+32\varepsilon/13}+\left(p^{3/26-46\varepsilon/13}\right)^{13/46}p^{89/92}\ll p^{1-\varepsilon}$
@@
$S_{p}(f)\ll p^{16/23}+p^{89/92}\leq p^{89/92}=p^{1-3/92}.$
Hence, recalling the definition (2.1), we see that this establishes the basis of induction for Theorems 2.1 and 2.3.
@@ 3.3. Induction step
where $t_{r}(\varepsilon)$ is given by (2.3), to ensure that (3.5) holds, we derive $S_{p}(f)\ll p^{1-\varepsilon/t_{r}(\varepsilon)}=p^{1-\kappa_{r}(\varepsilon)}$, where $\kappa_{r}(\varepsilon)$ is as in (2.2).
@@
4. Proof of Theorem 2.5
4.1. Exponential sums modulo prime powers
We first recall the following bound on the number of zeros of sparse polynomials in ${\mathbb{F}}_{p}$, which follows instantly from [6, Lemma 7].
Lemma 4.1.
Let $p$ be any prime and let $f\in{\mathbb{F}}_{p}[X]$ be be as in (1.2) and (1.3) with $q=p$. Let
$D=\max_{1\leq i\neq j\leq r}\gcd(e_{i}-e_{j},p-1).$
Then, the equation
$f(X)=0,\qquad x\in{\mathbb{F}}_{p},$
has at most $O\left(p^{1-1/r}D^{1/r}\right)$ solutions.
We now estimate
Lemma 4.2.
Let $p$ be any prime and let $m\geq 2$ be any integer. Assume that $f$ is as in (1.2) satisfying (1.3) with $q=p$. Assume that for all $1\leq i\leq r$ we have
(4.1) $\gcd(e_{i},p^{m})\leq p^{m/3}.$
Then, we have
$S_{p^{m}}(f)\ll p^{m-1/r}D^{1/r}$
where $D$ is as in Lemma 4.1.
Proof.
We start as in the proof of [4, Proposition 8]. Let us first set,
$n=\left\lceil m/2\right\rceil.$
Writing
$S_{p^{m}}(f)=\frac{1}{p^{n}}\sum_{x\in{\mathbb{Z}}_{p^{m}}^{*}}\sum_{1\leq y\leq p^{n}}\mathbf{e}_{p^{m}}(f(x+p^{n}y)),$
and noting that
$f(x+p^{n}y)=f(x)+p^{n}f^{\prime}(x)y\pmod{p^{m}},$
we have
$|S_{p^{m}}(f)|$ $\leq\frac{1}{p^{n}}\sum_{x\in{\mathbb{Z}}_{p^{m}}^{*}}\left|\sum_{1\leq y<p^{n}}\mathbf{e}_{p^{m-n}}(f^{\prime}(x)y)\right|$
=#$=\#\left\{x\in{\mathbb{Z}}_{p^{m}}^{*}:\leavevmode\nobreak\ f^{\prime}(x)\equiv 0\pmod{p^{m-n}}\right\}.$
We now write
$f^{\prime}(x)=\sum_{i=1}^{r}a_{i}e_{i}x^{e_{i}-1}=p^{u}\sum_{i=1}^{r}b_{i}x^{e_{i}-1},$
with some integers $u$ and $b_{1},\ldots,b_{r}$ satisfying $\gcd(b_{1},\ldots,b_{r},p)=1$. Since $m\geq 2$, the condition (4.1) implies that $u\leq m/3<m-n$. In particular,
$f^{\prime}(x)\equiv 0\pmod{p^{m-n}}$
implies
$\sum_{i=1}^{r}b_{i}x^{e_{i}-1}\equiv 0\pmod{p}.$
[…]
4.2. Concluding the proof
If
$q=\prod_{p\mid q}p^{m_{p}}$
is the prime factorisation of $q$, then one can write
(4.2) $S_{q}(f)=\prod_{p\mid q}S_{p^{m_{p}}}(\lambda_{p}f),$
with some integers $\lambda_{p}$ such that $\gcd(\lambda_{p},p)=1$, see, for example, [21, Equation (12.21)].
We consider the following system of three independent conditions:
(4.3a) $\gcd(e_{j},p-1)\leq p^{1/4},\qquad 1\leq u\leq r,$
(4.3b) $\gcd(e_{i}-e_{j},p-1)\leq p^{3/5},\qquad 1\leq i\neq j\leq r,$
(4.3c) $\gcd(e_{i},e_{j},p-1)\leq p^{3/52},\qquad 1\leq i\neq j\leq r.$
Note that the condition (4.3b) implies that we have $D\leq p^{3/5}$ in Lemmas 4.1 and 4.2.
Now, we also set
$E=\prod_{1\leq i<j\leq r}(e_{i}-e_{j})$
and consider the sets of primes
${\mathcal{P}}_{1}=\left\{p:\leavevmode\nobreak\ \gcd\left(p^{m_{p}},e_{1}\ldots e_{r}\right)>p^{m_{p}/3}\right\},$
${\mathcal{P}}_{2}=\left\{p:\leavevmode\nobreak\ m_{p}>20r^{2}\text{ and }E\equiv 0\pmod{p^{\lfloor m_{p}/(10r)\rfloor}}\right\},$
${\mathcal{P}}_{3}=\bigl{\{}p\not\in{\mathcal{P}}_{1}:\leavevmode\nobreak\ m_{p}\leq 20r^{2}\text{ and }p\text{ does not satisfy}$
[…]
Let us define
$q_{i}=\prod_{p\in{\mathcal{P}}_{i}}p^{m_{p}},\qquad i=1,2,3,$
and follow the arguments as in the proof of in [4, Proposition 11].
Note that
$|E|\leq d^{r(r-1)/2}<d^{r^{2}/2}.$
Hence for $q_{1}$ and $q_{2}$ we have the estimates
(4.4) $q_{1}<d^{3r}\qquad\mbox{and}\qquad q_{2}<d^{20r^{3}},$
see [4, Equations (3.11) and (3.12)].
For any $p\in{\mathcal{P}}_{3}$, failing one of the conditions (4.3a), (4.3b) and (4.3c) means
$p^{1/4}\leq d,\quad\text{or}\quad p^{3/5}\leq d,\quad\text{or}\quad p^{3/52}\leq d.$
Hence, in any case we have $p\leq d^{52/3}$. Therefore, we trivially have
(4.5) $q_{3}<\prod_{p\leq d^{52/3}}p^{20r^{2}}.$
By the classical result of Rosser and Schoenfeld [31, Theorem 9], we see that (4.5) implies
(4.6) $q_{3}<\exp\left(1.02\cdot 20r^{2}d^{52/3}\right)\leq\exp\left(21r^{2}d^{52/3}\right)$
Next, taking ${\mathcal{P}}={\mathcal{P}}_{1}\cup{\mathcal{P}}_{2}\cup{\mathcal{P}}_{3}$, it follows from (4.2) that
(4.7) $S_{q}(f)\ll q_{1}q_{2}q_{3}\prod_{\begin{subarray}{c}p\not\in{\mathcal{P}}\\ p\mid q\end{subarray}}\left(\sum_{x\in{\mathbb{Z}}_{p^{m_{p}}}^{*}}\mathbf{e}_{p^{m_{p}}}(\lambda_{p}f(x))\right).$
We now follow the argument as in the proof of [4, Proposition 11], and decompose the product of the sums on the right above as the product of three sums according to,
$m_{p}=1,\qquad 1<m_{p}\leq 20r^{2},\qquad m_{p}>20r^{2},$
and denote these products $S_{1},S_{2}$, and $S_{3}$ respectively
To estimate $S_{1}$, instead of the conditions of Theorem 2.1 corresponding to $\varepsilon=3/184$ we impose more stringent conditions (4.3a), (4.3b) and (4.3c), which guarantee that the bound of Theorem 2.1 holds with $\kappa_{r}(3/184)\geq\sigma_{r}$, as defined in (2.5), and we derive
(4.8) $S_{1}\ll Q_{1}^{1-\sigma_{r}+o(1)},\quad\mathrm{where}\ Q_{1}=\prod_{\begin{subarray}{c}p\not\in{\mathcal{P}},\leavevmode\nobreak\ p\mid q\\ m_{p}=1\end{subarray}}p,$
where, here and below, the term $Q^{o(1)}$ from the classical bound on the divisor function [21, Equation (1.81)].
Moreover, applying Lemma 4.2 for $m_{p}\geq 2$, Since for $p\not\in{\mathcal{P}}$ with $m_{p}\geq 2$, the condition (4.3b) that we have $D\leq p^{3/5}$. Hence
(4.9) $S_{2}\ll Q_{2}^{1-1/(50r^{3})+o(1)},\quad\mathrm{where}\ Q_{2}=\prod_{\begin{subarray}{c}p\not\in{\mathcal{P}},\leavevmode\nobreak\ p\mid q\\ 2\leq m_{p}\leq 20r^{2}\end{subarray}}p^{m_{p}}.$
Finally to estimate $S_{3}$, since we run over $m_{p}>20r^{2}$ while $p\not\in{\mathcal{P}}_{2}$, we apply [4, Proposition 7], and get
(4.10) $|S_{3}|\leq Q_{3}^{1-1/(10r^{2})+o(1)},\quad\mathrm{where}\ Q_{3}=\prod_{\begin{subarray}{c}p\not\in{\mathcal{P}},\leavevmode\nobreak\ p\mid q\\ m_{p}>20r^{2}\end{subarray}}p^{m_{p}}.$
For any integer $r\geq 1$ we have
$\sigma_{r}\leq\frac{1}{50r^{3}}.$
Combining (4.7), (4.8), (4.9), and (4.10), we have
$S_{q}(f)\ll q_{1}q_{2}q_{3}\left(Q_{1}Q_{2}Q_{3}\right)^{1-\sigma_{r}+o(1)}=q^{1-\sigma_{r}+o(1)}\left(q_{1}q_{2}q_{3}\right)^{\sigma_{r}}.$
Combining (4.4) and (4.6), we see that
$q_{1}q_{2}q_{3}\leq d^{3r+20r^{3}}\exp\left(21r^{2}d^{52/3}\right)\leq\exp\left(r^{3}d^{18}\right),$
concluding the proof.